OpenStudy (kaylala):

proving identities: cos x (secx-cosx) = sin^2 x

OpenStudy (usukidoll):

http://www.twiddla.com/1469723

OpenStudy (anonymous):

cos x(secx-cosx)=sin^2 x cos x (1/cos x - cos x) = sin^2 x cos x (1-cos^2 x / cos)=sin^2 x then cancel cos x 1-cos^2 x = sin^ 2 x sin^2 x = sin^2 x

OpenStudy (usukidoll):

yay